OpenStudy (anonymous):

two crooks rob a bank and flee to the east at 66 mph. In 35 mintues, the police follow them into the helicopter, flying at 132 mph. How long will it take for the police to overtake the robbers?

OpenStudy (anonymous):

Car: 66(mph)/60(minutes in an hour) = 1.1 miles per minute 1.1 x 35 = 38.5miles Chopper: 132/60 = 2.2 miles per minute 38.5(miles)/2.2(miles per minute) = 17.5 minutes Now, the car has been travelling for another 17.5 minutes while the chopper has made up this distance, so we need to work out how much further it will have travelled: 17.5 x 1.1 = 19.25miles. So, 38.5 + 19.25 = 57.75 (total miles travelled by car) Back to the Chopper: 57.75/2.2 = 26.25minutes...... I think :)

OpenStudy (anonymous):

Wow that was very discriptive! :) Thankyou so much! I was not sure how to measure out the the times with the chopper and car.. Difficult lol :)